# Pre-training Cross-lingual Open Domain Question Answering with Large-scale Synthetic Supervision

### v1

@@ A.3.2 Multilingual Open Domain Question Answering
Rewrite sentences into short and precise questions, using given question words and answers: Sentence: 2月、竇憲は左校尉の耿夔を遣わし、金微山において北匈奴の単于を包囲しこれを大いに破り、単于の母の閼氏を捕虜とした。 Question word: 誰 Answer: 匈奴 Transformed Question: 2月に金微山で竇憲の遣わした左校尉の耿夔が包囲し大いに破ったのは誰の単于ですか? Sentence: この町を法人化する法はリチャード・キャズウェルが提出し、キャズウェルはここを本拠地とし、後の1776年から1780年までノースカロライナ州の初代知事となった。 Question word: どこ Answer: ノースカロライナ州 Transformed Question: リチャード・キャズウェルが初代知事となったのはどこですか? Sentence: これより以前、司空張華は司馬倫に疎まれて誅殺されていた。 Question word: 誰 Answer: 張華 Transformed Question: 誰がこれより以前に司馬倫に疎まれて誅殺されていたのですか? Sentence: 魯迅はこの無支祁が孫悟空の先祖・源流ではないかと推測した。 Question word: 誰 Answer: 魯迅 Transformed Question: 誰はこの無支祁が孫悟空の先祖・源流ではないかと推測したのでしょうか? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: In February, Dou Xian sent Zuo’s lieutenant, Geng Kui, to besiege and defeat the Northern Xiongnu Danyu at Jinweishan, and took Danyu’s mother, the Yan family, prisoner. Question word: Who Answer: Xiongnu Transformed Question: In February, in Jinweishan, which was the land of Danyu that was besieged and severely defeated by Geng Ku, the commander of the left school sent by Dou Xian? Sentence: The act to incorporate the town was introduced by Richard Caswell, who made it his home and later became North Carolina’s first governor from 1776 to 1780. Question word: Where Answer: North Carolina Transformed Question: Where did Richard Caswell become the first governor? Sentence: Before this, Zhang Hua was shunned by Sima Lun and killed. Question word: Who Answer: Zhang Hua Transformed Question: Who had been shunned and killed by Sima Lun before this? Sentence: Lu Xun surmised that this Mujiqi was the ancestor and origin of Sun Wukong. Question word: Who Answer: Lu Xun Transformed Question: Who could have guessed that Mujiqi was the ancestor/origin of Son Goku?
@@
Rewrite sentences into short and precise questions, using given question words and answers: Sentence: 전투에서 승리한 뒤, 오버워치는 10년간 계속해서 평화를 지켰으나 내분으로 인해 해산되었다. Question word: 누구 Answer: 오버워치 Transformed Question: 누구가 전투에서 승리한 뒤 10년 동안 평화를 지키다가 내분으로 인해 해산되었나요? Sentence: 그가 구단을 떠난 지 10년이 되는 2013년 4월, 스포르팅 리스본은 호날두를 100,000번째 회원으로 등록해 경의를 표했다. Question word: 누구 Answer: 스포르팅 리스본 Transformed Question: 누가 2013년 4월 그가 구단을 떠난 지 10년이 되는 해에 호날두를 100,000번째 회원으로 등록해 경의를 표했나요? Sentence: 19세기 후반에 아일랜드에는 독립과 토지개혁을 요구하는 운동이 크게 확산되었다. Question word: 어디 Answer: 아일랜드 Transformed Question: 19세기 후반에 독립과 토지개혁을 요구하는 운동이 크게 확산된 나라는 어디입니까? Sentence: 산탄젤로 다리 () 또는 하드리아누스의 다리는 로마에 있는 다리 가운데 하나이다. Question word: 어디 Answer: 로마 Transformed Question: 산탄젤로 다리가 있는 곳은 어디인가? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: After winning the battle, Overwatch continued to maintain peace for 10 years, but was disbanded due to internal strife. Question word: Who Answer: Overwatch Transformed Question: Who won the battle, kept the peace for ten years, and then disbanded due to infighting? Sentence: In April 2013, 10 years after he left the club, Sporting Lisbon paid tribute to Ronaldo by registering him as their 100,000th member. Question word: Who Answer: Sporting Lisbon Transformed Question: Who paid tribute to Ronaldo by registering him as their 100,000th member in April 2013, marking 10 years since he left the club? Sentence: In the late 19th century, movements calling for independence and land reform spread widely in Ireland. Question word: Where Answer: Ireland Transformed Question: In which country did the movement calling for independence and land reform spread significantly in the late 19th century? Sentence: Ponte Sant’Angelo () or Hadrian’s Bridge is one of the bridges in Rome. Question word: Where Answer: Rome Transformed Question: Where is the Ponte Sant’Angelo?
@@
Rewrite sentences into short and precise questions, using given question words and answers: Sentence: ولد كلاي في مقاطعة هانوفر بولاية فرجينيا في عام 1777، ولكنه انتقل إلى ليكسينغتون، كنتاكي في عام 1797. Question word: أين Answer: كنتاكي Transformed Question: أين انتقل ولد كلاي بعد ميلاده في مقاطعة هانوفر، ولاية فرجينيا في عام 1797؟ Sentence: وكان نظام شركة سي بي إس هو الأكثر تقدما وفاز بالمنافسة في كل مرة. Question word: من Answer: شركة سي بي إس Transformed Question: من كان لديه النظام الأكثر تقدما وفاز بالمنافسة في كل مرة؟ Sentence: كما الخطوط الجوية البريطانية تشغل أيضا صالة نادي التنفيذي بين البوابات B21 و B23. Question word: من Answer: الخطوط الجوية البريطانية Transformed Question: من يشغل صالة نادي التنفيذي بين البوابات B21 و B23؟ Sentence: وقد افتتح حديثاً في قصر العظم قاعتين تضمان بانوراما لحرب تشرين لتحريرية عام 1973، وبانوراما لحرب تموز 2006. Question word: أين Answer: قصر العظم Transformed Question: أين بانوراما لحرب تشرين لتحريرية عام 1973؟ Rewrite sentences into short and precise questions, using given question words and answers: Sentence: Clay was born in Hanover County, Virginia in 1777, but moved to Lexington, Kentucky in 1797. Question word: Where Answer: Kentucky Transformed Question: Where did Clay move after his birth in Hanover County, Virginia in 1797? Sentence: The CPS system was the most advanced and won the competition. Question word: Who Answer: CPS system Transformed Question: Who had the most advanced system and won the competition? Sentence: British Airways also operates the Teen Club lounge between gates B21 and B23. Question word: Who Answer: British Airways Transformed Question: Who operates the Executive Club lounge between gates B21 and B23? Sentence: Two halls were recently opened in Al-Azm Palace containing a panorama of the October Liberation War of 1973, and a panorama of the July War of 2006. Question word: Where Answer: Al-Azm Palace Transformed Question: Where is the panorama of the October Liberation War of 1973?
@@
Rewrite sentences into short and precise questions, using given question words and answers: Sentence: JatRapeth sbar Aaeg edRoupdii pRaN Haran. Question word: ek Answer: edRoupdii Transformed Question: JatRapeth sbar Aaeg ek pRaN Haran? Sentence: Apridek katarer rajdhanii edaHaet raishJar EkiT s/thaJii duutabas rJeech. Question word: ekathaJ Answer: katar Transformed Question: raishJar s/thaJii duutabasiT ekathaJ Abis/tht? Sentence: bharte Hajar Hajar manuSh AnaHaer mara JaJ, ikn/tu dhr/mpRcarkra taedr pRit Udasiin. Question word: ekathaJ Answer: bhart Transformed Question: ekathaJ Hajar Hajar manuSh AnaHaer mara JaJ EbNNG dhr/mpRcarkra taedr pRit Udasiin thaek? Sentence: EiT OJaishNNGTn -Er isJaTl-E Abis/tht ekhala jaJgaJ EkiT maechr bajar. Question word: ekathaJ Answer: isJaTl Transformed Question: EiT OJaishNNGTn ekathaJ ekhala jaJgaJ EkiT maechr bajar? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: Draupadi was the first to die on the journey. Question word: Who Answer: Draupadi Transformed Question: Who died first on the journey? Sentence: In addition, Russia has a permanent embassy in Doha, the capital of Qatar. Question word: Where Answer: Qatar Transformed Question: Where is the permanent embassy of Russia located? Sentence: Thousands of people die of starvation in India, but missionaries are indifferent to them. Question word: Where Answer: India Transformed Question: Where are thousands of people dying of starvation and the missionaries are indifferent to them? Sentence: It is an open-air fish market located in Seattle, Washington. Question word: Where Answer: Seattle Transformed Question: Where is an open air fish market in Washington?
@@
[Uncaptioned image] Rewrite sentences into short and precise questions, using given question words and answers: Sentence: The main crops in this village are rice, sugarcane, mango, groundnut, vegetables etc. Question word: Who Answer: mango Transformed Question: Which is one of the main crops in this village? Sentence: China accounted for a significant increase in world emissions during this period. Question word: Where Answer: China Transformed Question: Where in the world has caused the significant increase in emissions during this time? Sentence: Among these, the famous Sri Vasavi Kanyaka Parameshwari Devi Temple is located in the town of Penugonda in the West Godavari district of the state of Andhra Pradesh. Question word: Who Answer: Penugonda Transformed Question: Which town in West Godavari district of Andhra Pradesh state has the famous Sri Vasavi Kanyaka Parameshwari Devi temple? Sentence: Seeing Satyaki being stopped by Kritavarma, Drona went towards Dharmaraja. Question word: Who Answer: Dharmaraja Transformed Question: To whom did Drona go when he saw Kritavarma stopping Satyaki?

### v3

@@ C.2 Case Study
Prompt E.3: Japanese Example & Translation Rewrite sentences into short and precise questions, using given question words and answers:{CJK} UTF8min Sentence: 2月、竇憲は左校尉の耿夔を遣わし、金微山において北匈奴の単于を包囲しこれを大いに破り、単于の母の閼氏を捕虜とした。 Question word: 誰 Answer: 匈奴 Transformed Question: 2月に金微山で竇憲の遣わした左校尉の耿夔が包囲し大いに破ったのは誰の単于ですか? Sentence: この町を法人化する法はリチャード・キャズウェルが提出し、キャズウェルはここを本拠地とし、後の1776年から1780年までノースカロライナ州の初代知事となった。 Question word: どこ Answer: ノースカロライナ州 Transformed Question: リチャード・キャズウェルが初代知事となったのはどこですか? Sentence: これより以前、司空張華は司馬倫に疎まれて誅殺されていた。 Question word: 誰 Answer: 張華 Transformed Question: 誰がこれより以前に司馬倫に疎まれて誅殺されていたのですか? Sentence: 魯迅はこの無支祁が孫悟空の先祖・源流ではないかと推測した。 Question word: 誰 Answer: 魯迅 Transformed Question: 誰はこの無支祁が孫悟空の先祖・源流ではないかと推測したのでしょうか? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: In February, Dou Xian sent Zuo’s lieutenant, Geng Kui, to besiege and defeat the Northern Xiongnu Danyu at Jinweishan, and took Danyu’s mother, the Yan family, prisoner. Question word: Who Answer: Xiongnu Transformed Question: In February, in Jinweishan, which was the land of Danyu that was besieged and severely defeated by Geng Ku, the commander of the left school sent by Dou Xian? Sentence: The act to incorporate the town was introduced by Richard Caswell, who made it his home and later became North Carolina’s first governor from 1776 to 1780. Question word: Where Answer: North Carolina Transformed Question: Where did Richard Caswell become the first governor? Sentence: Before this, Zhang Hua was shunned by Sima Lun and killed. Question word: Who Answer: Zhang Hua Transformed Question: Who had been shunned and killed by Sima Lun before this? Sentence: Lu Xun surmised that this Mujiqi was the ancestor and origin of Sun Wukong. Question word: Who Answer: Lu Xun Transformed Question: Who could have guessed that Mujiqi was the ancestor/origin of Son Goku?
@@
Prompt E.4: Korean Example & Translation Rewrite sentences into short and precise questions, using given question words and answers:{CJK} UTF8mj Sentence: 전투에서 승리한 뒤, 오버워치는 10년간 계속해서 평화를 지켰으나 내분으로 인해 해산되었다. Question word: 누구 Answer: 오버워치 Transformed Question: 누구가 전투에서 승리한 뒤 10년 동안 평화를 지키다가 내분으로 인해 해산되었나요? Sentence: 그가 구단을 떠난 지 10년이 되는 2013년 4월, 스포르팅 리스본은 호날두를 100,000번째 회원으로 등록해 경의를 표했다. Question word: 누구 Answer: 스포르팅 리스본 Transformed Question: 누가 2013년 4월 그가 구단을 떠난 지 10년이 되는 해에 호날두를 100,000번째 회원으로 등록해 경의를 표했나요? Sentence: 19세기 후반에 아일랜드에는 독립과 토지개혁을 요구하는 운동이 크게 확산되었다. Question word: 어디 Answer: 아일랜드 Transformed Question: 19세기 후반에 독립과 토지개혁을 요구하는 운동이 크게 확산된 나라는 어디입니까? Sentence: 산탄젤로 다리 () 또는 하드리아누스의 다리는 로마에 있는 다리 가운데 하나이다. Question word: 어디 Answer: 로마 Transformed Question: 산탄젤로 다리가 있는 곳은 어디인가? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: After winning the battle, Overwatch continued to maintain peace for 10 years, but was disbanded due to internal strife. Question word: Who Answer: Overwatch Transformed Question: Who won the battle, kept the peace for ten years, and then disbanded due to infighting? Sentence: In April 2013, 10 years after he left the club, Sporting Lisbon paid tribute to Ronaldo by registering him as their 100,000th member. Question word: Who Answer: Sporting Lisbon Transformed Question: Who paid tribute to Ronaldo by registering him as their 100,000th member in April 2013, marking 10 years since he left the club? Sentence: In the late 19th century, movements calling for independence and land reform spread widely in Ireland. Question word: Where Answer: Ireland Transformed Question: In which country did the movement calling for independence and land reform spread significantly in the late 19th century? Sentence: Ponte Sant’Angelo () or Hadrian’s Bridge is one of the bridges in Rome. Question word: Where Answer: Rome Transformed Question: Where is the Ponte Sant’Angelo?
@@
Prompt E.5: Arabic Example & Translation Rewrite sentences into short and precise questions, using given question words and answers: Sentence: ولد كلاي في مقاطعة هانوفر بولاية فرجينيا في عام 1777، ولكنه انتقل إلى ليكسينغتون، كنتاكي في عام 1797. Question word: أين Answer: كنتاكي Transformed Question: أين انتقل ولد كلاي بعد ميلاده في مقاطعة هانوفر، ولاية فرجينيا في عام 1797؟ Sentence: وكان نظام شركة سي بي إس هو الأكثر تقدما وفاز بالمنافسة في كل مرة. Question word: من Answer: شركة سي بي إس Transformed Question: من كان لديه النظام الأكثر تقدما وفاز بالمنافسة في كل مرة؟ Sentence: كما الخطوط الجوية البريطانية تشغل أيضا صالة نادي التنفيذي بين البوابات B21 و B23. Question word: من Answer: الخطوط الجوية البريطانية Transformed Question: من يشغل صالة نادي التنفيذي بين البوابات B21 و B23؟ Sentence: وقد افتتح حديثاً في قصر العظم قاعتين تضمان بانوراما لحرب تشرين لتحريرية عام 1973، وبانوراما لحرب تموز 2006. Question word: أين Answer: قصر العظم Transformed Question: أين بانوراما لحرب تشرين لتحريرية عام 1973؟ Rewrite sentences into short and precise questions, using given question words and answers: Sentence: Clay was born in Hanover County, Virginia in 1777, but moved to Lexington, Kentucky in 1797. Question word: Where Answer: Kentucky Transformed Question: Where did Clay move after his birth in Hanover County, Virginia in 1797? Sentence: The CPS system was the most advanced and won the competition. Question word: Who Answer: CPS system Transformed Question: Who had the most advanced system and won the competition? Sentence: British Airways also operates the Teen Club lounge between gates B21 and B23. Question word: Who Answer: British Airways Transformed Question: Who operates the Executive Club lounge between gates B21 and B23? Sentence: Two halls were recently opened in Al-Azm Palace containing a panorama of the October Liberation War of 1973, and a panorama of the July War of 2006. Question word: Where Answer: Al-Azm Palace Transformed Question: Where is the panorama of the October Liberation War of 1973?
@@
Prompt E.6: Bengali Example & Translation Rewrite sentences into short and precise questions, using given question words and answers: Sentence: JatRapeth sbar Aaeg edRoupdii pRaN Haran. Question word: ek Answer: edRoupdii Transformed Question: JatRapeth sbar Aaeg ek pRaN Haran? Sentence: Apridek katarer rajdhanii edaHaet raishJar EkiT s/thaJii duutabas rJeech. Question word: ekathaJ Answer: katar Transformed Question: raishJar s/thaJii duutabasiT ekathaJ Abis/tht? Sentence: bharte Hajar Hajar manuSh AnaHaer mara JaJ, ikn/tu dhr/mpRcarkra taedr pRit Udasiin. Question word: ekathaJ Answer: bhart Transformed Question: ekathaJ Hajar Hajar manuSh AnaHaer mara JaJ EbNNG dhr/mpRcarkra taedr pRit Udasiin thaek? Sentence: EiT OJaishNNGTn -Er isJaTl-E Abis/tht ekhala jaJgaJ EkiT maechr bajar. Question word: ekathaJ Answer: isJaTl Transformed Question: EiT OJaishNNGTn ekathaJ ekhala jaJgaJ EkiT maechr bajar? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: Draupadi was the first to die on the journey. Question word: Who Answer: Draupadi Transformed Question: Who died first on the journey? Sentence: In addition, Russia has a permanent embassy in Doha, the capital of Qatar. Question word: Where Answer: Qatar Transformed Question: Where is the permanent embassy of Russia located? Sentence: Thousands of people die of starvation in India, but missionaries are indifferent to them. Question word: Where Answer: India Transformed Question: Where are thousands of people dying of starvation and the missionaries are indifferent to them? Sentence: It is an open-air fish market located in Seattle, Washington. Question word: Where Answer: Seattle Transformed Question: Where is an open air fish market in Washington?
@@
Prompt E.7: Telugu Example & Translation [Uncaptioned image] Rewrite sentences into short and precise questions, using given question words and answers: Sentence: The main crops in this village are rice, sugarcane, mango, groundnut, vegetables etc. Question word: Who Answer: mango Transformed Question: Which is one of the main crops in this village? Sentence: China accounted for a significant increase in world emissions during this period. Question word: Where Answer: China Transformed Question: Where in the world has caused the significant increase in emissions during this time? Sentence: Among these, the famous Sri Vasavi Kanyaka Parameshwari Devi Temple is located in the town of Penugonda in the West Godavari district of the state of Andhra Pradesh. Question word: Who Answer: Penugonda Transformed Question: Which town in West Godavari district of Andhra Pradesh state has the famous Sri Vasavi Kanyaka Parameshwari Devi temple? Sentence: Seeing Satyaki being stopped by Kritavarma, Drona went towards Dharmaraja. Question word: Who Answer: Dharmaraja Transformed Question: To whom did Drona go when he saw Kritavarma stopping Satyaki?